# Building An Open Architecture for AI-RAN Convergence in 6G

### v1

@@ IV Converged AI-RAN and O-RAN Architecture
Building on the overall network architecture shown in Fig. 1 and the requirements identified above, Figs. 3 and  4 detail the specific AI-RAN extensions to O-RAN components and their interfaces. This converged architecture has been designed to enable two concurrent workflows for the orchestration, deployment, and execution of AI workloads, i.e., batch and real-time. In batch mode, AI tasks are submitted to the centralized orchestrator, and distributed when/where compute capacity is available. This can be used, for example, for AI training, or general AI processing that requires acceleration on GPU but does not have constraints on the deployment location and timing of the deployment. In real-time mode, AI developers submit jobs for direct execution to a specific cluster (or geographical area of interest) and the request will be accommodated by the orchestrator as long as resources are available.
@@
Figure 3: Extension of the O-RAN SMO to support AI-RAN use cases and orchestration. The components in orange, together with the SMOS, are part of the O-RAN SMO architecture. Components in blue represent extensions to accommodate AI-RAN requirements.
@@
Figure 3 illustrates the architectural extension of the SMO toward an AI-enabled SMO (AI-SMO). Besides O-RAN SMO components connected through the SMOS bus, a set of new micro-services covers the following functionalities: resource allocation, AI workload automation, authentication and workload validation for the AI components, and exposure of interfaces to interact with the AI-RAN elements. By embedding the AI-RAN orchestrator in the SMO, we provide the orchestrator with (i) access to granular monitoring data on the infrastructure already collected through the O-RAN O1 and O2 interfaces; and (ii) a global, centralized viewpoint on the distributed RAN infrastructure that enables a holistic approach to allocation of resources to either AI or RAN workloads. The workload validation and AI user authentication address security concerns through an authorization process that vets users and AI tasks before they are admitted to the system. It validates the identity of AI users through the northbound interface to handle requests for training or inference execution. Upon successful validation, the component issues an authorization token granting access to the appropriate resources. Then, upon availability of resources determined by the orchestrator, AI task requests are forwarded to the AI workloads automation component, which triggers their deployment to the AI-RAN infrastructure sites. This components operates in parallel to equivalent services for RAN workloads, including the Network Function Orchestration (NFO) and Federated O-Cloud Orchestration and Management (FOCOM) services, with which it coordinates through the SMOS. Batch AI and RAN workloads are deployed to the selected AI-RAN edge sites using an extension of the O2 interface (AI-O2), which also communicates to the edge sites that resource allocation schedules or policies defined by the AI-RAN Orchestration service. Quality of Service (QoS) policies for AI-O2 traffic can further help by prioritizing RAN workloads over less time-sensitive AI tasks.
@@
Figure 4: Architecture for an edge cloud site that supports AI-RAN solutions on a compute cluster managed by the AI-O-Cloud. The top part of the figure shows the AI workloads and RAN workloads.
@@
At the other end of the AI-O2 interface, the AI-RAN edge site, shown in Fig. 4, is responsible for the granular, container-level service orchestration. Here, the O-Cloud extends into an AI-ready version that can manage containerized workloads for AI and RAN solutions. Figure 4 also shows the taxonomy of workloads that can be executed on AI-RAN sites. These include the AI-driven Central Units (CUs) and Distributed Units (DUs), with AI-for-RAN solutions (e.g., dApps [14]) providing intelligence within the protocol stack. An additional AI-for-RAN category, outside of the RAN CU/DUs, is represented by AI-accelerated Near-RT RICs, which can be hosted at the edge to satisfy near-real-time inference and control constraints. AI-on-RAN solutions can include training for AI models, generative AI services (e.g., chatbots), sensing based on AI, and other edge inference tasks, among others. As an instance, Fig. 5 illustrates the GPU utilization for two GPU-accelerated nodes in a prototype of an AI-RAN cluster, based on [15]. The system shares two applications, a GPU-based DU (based on NVIDIA ARC), with 15 connected users exchanging up to 1.5 Gbps of downlink traffic, and a chatbot using an Ollama-based LLM. The first represents an AI-for-RAN workload, and the second is an AI-on-RAN task. In this instance, the system has enough GPU capacity to support both solutions, showing the potential of RAN infrastructure sharing.
Figure 5: GPU utilization over time for a cluster with 2 GPUs (maximum GPU utilization at 200%), with AI-for-RAN (GPU-based DU) and AI-on-RAN (LLM chatbot) dynamic workloads.
@@
The AI-RAN site is equipped with compute (including AI accelerators, e.g., GPUs), storage, and networking resources that are presented and organized through a container orchestration platform (e.g., Kubernetes or RedHat OpenShift). Both of these components are managed by the AI-O-Cloud Infrastructure Management Service (IMS) and Deployment Management Service (DMS). These extend O-RAN services that terminate the O2 interface for infrastructure configuration and workload lifecycle management, respectively. In the AI-O-Cloud, they also translate policies and schedules determined by the SMO into AI infrastructure configurations and AI workload deployment, while enabling coexistence with the RAN.
@@
In Fig. 6, we propose a procedure for the real-time deployment of AI workloads that combines (i) pre-authentication with the AI-SMO, with an indication of computing requirements at specific AI-RAN sites; (ii) the definition of a resource sharing policy in the AI-RAN Orchestrator, which is then communicated to the edge site through AI-O2; (iii) the use of a Multi-access Edge Computing (MEC) interface, adopting the ETSI standards for edge computing, by the pre-authenticated users to submit real-time AI workloads to a specific AI-RAN site; and (iv) involvement of the DMS, which allocates resources and deploys the AI services with minimal latency, as long as the amount of resources required is compatible with the policy defined by the AI-RAN Orchestrator. If not, the AI task is not admitted, and the process resumes at the AI-SMO, where the user can explore options to increase priority of the task (e.g., through different monetization tiers) or submit it as a batch workload or to a different AI-RAN site.
@@ V Conclusions
We presented an end-to-end network architecture for AI-RAN coexistence. After reviewing the AI-RAN Alliance activities and O-RAN architecture, we discussed the challenges associated to operating AI and RAN on the same infrastructure. We then proposed an extension of the O-RAN architecture and O-RAN components that enables the deployment and management of AI workloads on the shared RAN infrastructure, including different modes for managing AI workloads with different requirements for deployment latency and location.

### v2

@@ IV Converged AI-RAN and O-RAN Architecture
Building on the architecture in Fig. 1 and the requirements identified above, Figs. 2 and 3 detail the specific AI-RAN extensions to O-RAN components and their interfaces. This converged architecture has been designed to enable two concurrent workflows for the orchestration, deployment, and execution of AI workloads, i.e., batch and real-time. In batch mode, AI tasks are submitted to the centralized orchestrator, and distributed when/where compute capacity is available. This can be used, for example, for AI training, or general AI processing that requires acceleration on GPU but does not have constraints on the deployment location and timing of the deployment. In real-time mode, AI developers submit jobs for direct execution to a specific cluster (or geographical area of interest) and the request will be accommodated by the orchestrator according to resource availability.
@@
Figure 2: Extension of the O-RAN SMO to support AI-RAN orchestration. The components in light blue, together with the SMOS, are part of the O-RAN SMO architecture. Components in dark blue represent extensions to accommodate AI-RAN requirements.
IV-A AI-SMO and AI-RAN Site
@@
AI-SMO. Figure 2 illustrates the architectural extension of the SMO toward an AI-enabled SMO (AI-SMO). Besides O-RAN SMO components connected through the SMOS bus, a set of new micro-services covers the following functionalities: resource allocation, AI workload automation, authentication and workload validation for the AI components, and exposure of interfaces to interact with the AI-RAN elements. By embedding the AI-RAN orchestrator in the SMO, we provide the orchestrator with (i) access to granular monitoring data on the infrastructure already collected through the O-RAN O1 and O2 interfaces; and (ii) a global, centralized viewpoint on the distributed RAN infrastructure that enables a holistic approach to allocation of resources to either AI or RAN workloads. The workload validation and AI user authentication provide access control with an authorization process that vets users and AI tasks before they are admitted to the system. It validates the identity of AI users through the northbound interface to handle requests for training or inference execution. Upon successful validation, the service issues an authorization token granting access to the appropriate resources. The authentication mechanism can be tailored to the specific deployment operational constraints, thanks to the disaggregated design of the AI-SMO. Then, upon availability of resources determined by the orchestrator, AI task requests are forwarded to the AI workloads automation component, which triggers their deployment to the AI-RAN infrastructure sites. This component operates in parallel to equivalent services for RAN workloads, including the Network Function Orchestration (NFO) and Federated O-Cloud Orchestration and Management (FOCOM) services, with which it coordinates through the SMOS. Batch AI and RAN workloads are deployed to the selected AI-RAN edge sites using an extension of the O2 interface (AI-O2), which also communicates to the edge sites that resource allocation schedules or policies defined by the AI-RAN Orchestration service. Quality of Service (QoS) policies for AI-O2 traffic can further help by prioritizing RAN workloads over less time-sensitive AI tasks. Finally, operators can dynamically express geographically and temporally customized policies that determine AI and RAN sharing strategies.
@@
Figure 3: Architecture for an edge cloud site that supports AI-RAN solutions on a compute cluster managed by the AI-O-Cloud. The figure shows the AI workloads and RAN workloads (top), the orchestration components (middle), and the infrastructure and interfaces (bottom).
@@
AI-RAN Site. At the other end of the AI-O2 interface, the AI-RAN edge site, shown in Fig. 3, is responsible for the granular, container-level service orchestration. Here, the O-Cloud extends into an AI-ready version that can manage containerized workloads for AI and RAN solutions, and also potentially distributed core components such as a local User Plane Function (UPF). Figure 3 also shows the taxonomy of workloads that can be executed on AI-RAN sites. Besides AI-driven Central Units (CUs) and Distributed Units (DUs), AI-for-RAN solutions include, for example, dApps [14], real-time programmable components that extend the O-RAN architecture to provide intelligence within the protocol stack. An additional AI-for-RAN category, outside of the RAN CUs/DUs, is represented by AI-accelerated Near-RT RICs, which can be hosted at the edge to satisfy near-real-time inference and control constraints. AI-on-RAN solutions can include training for AI models, generative AI services (e.g., chatbots), sensing based on AI, and other edge inference tasks, among others.
@@
The AI-RAN site is equipped with compute (including AI accelerators, e.g., GPUs), storage, and networking resources. These are managed by a container orchestration platform (e.g., Kubernetes or Red Hat OpenShift), which also provides tenant’s workload isolation via namespaces. Both of these components are managed by the AI-O-Cloud Infrastructure Management Service (IMS) and Deployment Management Service (DMS). These extend O-RAN services that terminate the O2 interface for infrastructure configuration and workload lifecycle management, respectively. In the AI-O-Cloud, they also translate policies and schedules determined by the SMO into AI infrastructure configurations and AI workload deployment, while enabling coexistence with the RAN.
IV-B Scalability and Real-time Deployment Workflow
@@
The service-based nature of the AI-SMO favors scalability, with different components that can be scaled to handle increasing workloads submissions or large number of sites. At the same time, the AI batch submission workflow and specific instances of optimization for the resource allocation may introduce delays when applied to large-scale systems. Further, specific AI-on-RAN task may require low-latency deployments. Examples include AI inference services that are dynamically offloaded by or need to follow highly mobile users in a network, e.g., connected autonomous vehicles. Public safety sensing applications may need rapid deployment in response to an emergency. For this reason, in Fig. 4, we propose a procedure for the real-time deployment of AI workloads. It combines (i) pre-authentication with the AI-SMO, with an indication of computing requirements at specific AI-RAN sites; (ii) the definition of a resource sharing policy in the AI-RAN Orchestrator, which is then communicated to the edge site through AI-O2; (iii) the use of a Multi-access Edge Computing (MEC) interface, adopting the ETSI standards for edge computing, by the pre-authenticated users to submit real-time AI workloads to a specific AI-RAN site; and (iv) involvement of the DMS, which allocates resources and deploys the AI services with minimal latency, as long as the amount of resources required is compatible with the policy defined by the AI-RAN Orchestrator. If not, the AI task is not admitted, and the process resumes at the AI-SMO, where the user can explore options to increase priority of the task (e.g., through different monetization tiers) or submit it as a batch workload or to a different AI-RAN site.
IV-C Profiling the AI-RAN Site
We profile the feasibility of the approach with a prototype of the AI-RAN site, focusing on AI and RAN coexistence and on their deployment latency. The results are based on multiple, independent experiments executed on the Northeastern University X5G testbed [15], which implements a private 5G network with multiple commercial O-RAN Radio Units (RUs) and fronthaul infrastructure. We integrate in X5G a single-node OpenShift cluster with a Gigabyte server with an x86 CPU and an NVIDIA A100 GPU. For the RAN software, we use the NVIDIA Aerial GPU-accelerated physical layer and OpenAirInterface. For AI workloads, we use different Ollama-based LLMs (processing prompts), and a Convolutional Neural Network (CNN) or a ResNet model trained live on CIFAR-100. The AI-SMO configures the AI-RAN site GPU with Multi-Instance GPU (MIG), an NVIDIA technology for resource partitioning. For these experiments, the GPU is partitioned in 3, with 40 GB for RAN, 20 GB for LLM, and 10 GB for ResNet/CNN.
CNNResNetCNNResNet0202040406060—— Llama-3.2-3B —— —— Llama-3-8B —— MAC Throughput [Mbps]DL Th / RAN-onlyDL Th / AI-RAN Coexistence05510101515CRC Error Rate [error/s]CRC Error Rate / RAN-onlyCRC Error Rate / AI-RAN Coexistence
Figure 5: Median values for throughput and CRC error rate with RAN only (solid bars) or AI-RAN coexistence with LLM prompts and CNN/ResNet (dotted bars).
Figure 5 reports the median throughput and CRC error rate at the MAC layer, thus including retransmissions, with 95% confidence intervals computed with the bootstrap method. The solid bars report the performance without AI; dotted bars are for experiments with User Equipment (UE) traffic, the CNN/ResNet training, and the LLM being prompted. The performance is consistent across all experiments, validating AI-RAN coexistence even in a single-node site.
Figure 6 illustrates the sequential deployment latency for different LLMs models and parameter sizes. We profile the workload deployment latency from when the AI-RAN site receives the request to when the service is available, assuming a sequential deployment for the LLM, followed by the RAN and ResNet. As can be seen, the RAN and ResNet are deployed in 1-4 s, with the longer time following the deployment of the largest Llama model that fits the MIG partition. LLM instances require between 5.8-34.4 s (average). This shows that complex workloads can be managed on the same infrastructure in a matter of seconds, and provides an indication of the dynamicity with which such workloads can be (re)deployed by the AI-SMO.
0551010151520202525303035354040Qwen2.5-1.5BLlama-3.2-1BLlama-3.2-3BLlama-3-8BTime [s]LLMRANCNN
Figure 6: Workload deployment latency on the AI-RAN site.
@@ V Conclusions
We presented an end-to-end network architecture for AI-RAN coexistence. After reviewing the AI-RAN Alliance activities and O-RAN architecture, we discussed the challenges associated to operating AI and RAN on the same infrastructure. We then proposed an extension of the O-RAN architecture and O-RAN components that enables the deployment and management of AI workloads on the shared RAN infrastructure, including different modes for managing AI workloads with different requirements for deployment latency and location. Near-term extensions of this work include further development of the architecture into an end-to-end prototype serving as a platform to validate resource allocation strategies. In the medium term, it will also facilitate studying how interfaces and design can be streamlined into O-RAN and AI-RAN Alliances activities and specifications.